Express 98 as a sum or difference of two numbers whose powers are easier to
calculate and then use binomial theorem to compute 98 raised to the power 5.

Found 2 solutions by richard1234, chessace:
Answer by richard1234(7193)



You can finish the calculation.

Answer by chessace(471)

98^5 = (100-2)^5
= 10^10 - 5*2*10^8 +10*4*10^6 - 10*8*10^4 + 5*16*100 - 32
=10040008000
-01000800032
=09039207968
Not as fast as a modern calculator with enough digits, but most of them do not have enough digits.